# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ Appendix A Middle Convolution Tables For Nice 3-Tuples
$\frac{38}{60}$ $-z^{8}$ $\begin{pmatrix}z^{24}-z^{20}&\left(z^{28}+2z^{4}\right)u\>+2\left(z^{8}-z^{2}\right)\\ z^{14}+u-z^{2}-1&-z^{24}-u+z^{2}\end{pmatrix}$ $\begin{pmatrix}z^{8}&-z^{20}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -2z^{18}-z^{6}&z^{8}\end{pmatrix}$ $(z^{26},z^{26},z^{26},z^{42})$ 60 0
$\frac{50}{60}$ $-z^{20}$ $\begin{pmatrix}-u-z^{20}&u\\ -u-z^{20}-1&u+1\end{pmatrix}$ $\begin{pmatrix}-z^{20}&z^{20}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{24}+z^{6}+1&-z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},1)$ 60 360
@@
$\frac{35}{60}$ $-z^{5}$ $\frac{1}{3}\begin{pmatrix}-z^{5}f_{1}-2z^{10}+1&f_{1}-2z^{15}+3z^{10}+z^{5}\\ f_{1}+g_{1}&z^{5}f_{1}+2z^{10}+3z^{5}+2\end{pmatrix}$ $\begin{pmatrix}z^{5}&-z^{20}f_{1}-u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{25}u-z^{15}&z^{5}\end{pmatrix}$ $(q^{55},q^{55},q^{55},q^{75})$ 96 0
$\frac{50}{60}$ $-z^{20}$ $\begin{pmatrix}f_{1}-u-1+2z^{10}&-f_{1}+2u+1\\ f_{1}-u+2z^{10}&-f_{1}+u-z^{20}\end{pmatrix}$ $\begin{pmatrix}z^{20}&-f_{1}+2u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ f_{1}-u+1&z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},1)$ 96 360
@@
$\frac{34}{60}$ $-z^{4}$ $\begin{pmatrix}-z^{8}-z^{6}&z^{18}f_{1}+2z^{12}+4\\ z^{14}-z^{50}v+2z^{10}+u&u-z^{50}v+1\end{pmatrix}$ $\begin{pmatrix}z^{4}&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ \frac{1}{2}\left(z^{4}f_{1}+3z^{24}-z^{12}+z^{6}\right)&z^{4}\end{pmatrix}$ $(z^{28},z^{28},z^{28},z^{36})$ 60 0
$\frac{46}{60}$ $-z^{16}$ $\begin{pmatrix}z^{14}-z^{4}+z^{2}&f_{2}\\ -v+z^{28}v+z^{10}-z^{8}&\left(z^{50}+z^{28}\right)v-u\end{pmatrix}$ $\begin{pmatrix}z^{16}&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ \frac{1}{2}\left(z^{22}+2z^{16}u+z^{10}-3z^{6}\right)&z^{16}\end{pmatrix}$ $(z^{22},z^{22},z^{22},z^{54})$ 60 0
@@ Appendix B Middle Convolution Tables of Nice 4-Tuples
A $\frac{1}{2}$ $-1$ $\begin{pmatrix}u&-u+1\\ u-1&-u+2\end{pmatrix}$ $\begin{pmatrix}0&1\\ -1&2\end{pmatrix}$ $\begin{pmatrix}1&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&1\end{pmatrix}$ $(1,1,1,1,1)$ 0
[…]
tableHere $q=\zeta_{5}$ and $u=q^{3}+q^{2}$
@@
[…]
tableHere $z=\zeta_{3}$ and $u=\zeta_{12}$
@@
A $\frac{5}{6}$ $-z^{5}$ $\begin{pmatrix}-f_{2}+1&\frac{1}{4}g_{1}\\ g_{2}&z^{5\>}+f_{2}\end{pmatrix}$ $\begin{pmatrix}f_{1}+2&\frac{1}{2}f_{2}\\ g_{3}&g_{4}\end{pmatrix}$ $\begin{pmatrix}z^{5}&\frac{1}{4}g_{5}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ g_{6}&z^{5}\end{pmatrix}$ $(z^{5},z^{5},z^{5},z^{5},z^{10})$ 360
@@
$\frac{5}{6}$ $-z^{4}+1$ $\begin{pmatrix}1&1\\ 0&z^{8}\end{pmatrix}$ $\begin{pmatrix}z^{4}+1&z^{4}\\ -2&-1\end{pmatrix}$ $\begin{pmatrix}z^{8}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 2z^{8}&z^{8}\end{pmatrix}$ $(z^{8},z^{8},z^{8},z^{8},-z^{4})$ 45 72
@@
$\frac{7}{10}$ $-z^{12}$ $\begin{pmatrix}z^{12}-v&-z^{24}\\ -z^{18}+v&z^{6}\end{pmatrix}$ $\begin{pmatrix}-z^{24}+z^{12}&z^{12}\\ z^{24}-z^{6}v&z^{24}+1\end{pmatrix}$ $\begin{pmatrix}z^{12}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{24}+z^{12}-1&z^{12}\end{pmatrix}$ $(z^{24},z^{24},z^{24},z^{24},z^{24})$ 50 600
@@
$\frac{1}{6}$ $z^{10}$ $\begin{pmatrix}-z^{20}-z^{16}&-z^{16}-z^{10}\\ z^{16}-1&z^{16}\end{pmatrix}$ $\begin{pmatrix}-z^{10}&0\\ z^{10}&1\end{pmatrix}$ $\begin{pmatrix}-z^{10}&u-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{10}&-z^{10}\end{pmatrix}$ $-(z^{10},z^{10},z^{10},z^{10},-z^{20})$ 90 360
$\frac{5}{6}$ $-z^{10}+1$ $\begin{pmatrix}y+z^{20}&y+z^{10}-2\\ -y&-y+1\end{pmatrix}$ $\begin{pmatrix}z^{20}&0\\ -z^{20}&1\end{pmatrix}$ $\begin{pmatrix}z^{20}&u-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{20}&z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},z^{20},-z^{10})$ 90 360
@@
$\frac{9}{10}$ $-z^{24}$ $\begin{pmatrix}1&-u-1\\ 0&z^{24}\end{pmatrix}$ $\begin{pmatrix}z^{24}+1&z^{6}v\\ u&0\end{pmatrix}$ $\begin{pmatrix}z^{24}&u+1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ u+z^{12}&z^{24}\end{pmatrix}$ $(z^{18},z^{18},z^{18},z^{18},z^{48})$ 50 600
@@
$\frac{5}{6}$ $-z^{2}+1$ $\begin{pmatrix}0&-1\\ -1&0\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{2}\end{pmatrix}$ $\begin{pmatrix}z^{2}&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{2}\end{pmatrix}$ $(-z^{3},z^{5},z^{5},z^{5},1)$ 40 72
@@
$\frac{5}{6}$ $-z^{2}+1$ $\begin{pmatrix}1&-w\\ 0&-1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-w^{2}\end{pmatrix}$ $\begin{pmatrix}-1&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-w\end{pmatrix}$ $(z^{3},z^{5},z^{3},z,1)$ 0
@@
$\frac{7}{12}$ $-z$ $\begin{pmatrix}z^{3}&0\\ -z^{11}&1\end{pmatrix}$ $\begin{pmatrix}1&z^{4}\\ 0&z^{3}\end{pmatrix}$ $\begin{pmatrix}z^{3}&-z^{4}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{11}&z^{3}\end{pmatrix}$ $(q^{9},q^{9},q^{9},q^{9},-1)$ 20 96
@@ Appendix C Middle Convolution Tables for Nice 5-Tuples
A $\frac{5}{6}$ $-w$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}0&w+1\\ -w&0\end{pmatrix}$ $\begin{pmatrix}w+1&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $(z^{11},z^{11},-z^{3},z^{11},z^{11},z^{7})$ 72
B $\frac{5}{6}$ $-w$ $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}0&-w-1\\ w&0\end{pmatrix}$ $\begin{pmatrix}w+1&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $(z^{3},z^{11},z^{3},z^{11},z^{11},-z^{3})$ 72

### v3

@@ Appendix A Middle Convolution Tables For Nice 3-Tuples
$\frac{38}{60}$ $z^{8}$ $\begin{pmatrix}z^{24}-z^{20}&\left(z^{28}+2z^{4}\right)u\>+2\left(z^{8}-z^{2}\right)\\ z^{14}+u-z^{2}-1&-z^{24}-u+z^{2}\end{pmatrix}$ $\begin{pmatrix}z^{8}&-z^{20}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -2z^{18}-z^{6}&z^{8}\end{pmatrix}$ $(z^{26},z^{26},z^{26},z^{42})$ 60 0
$\frac{50}{60}$ $z^{20}$ $\begin{pmatrix}-u-z^{20}&u\\ -u-z^{20}-1&u+1\end{pmatrix}$ $\begin{pmatrix}-z^{20}&z^{20}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{24}+z^{6}+1&-z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},1)$ 60 360
@@
$\frac{35}{60}$ $z^{5}$ $\frac{1}{3}\begin{pmatrix}-z^{5}f_{1}-2z^{10}+1&f_{1}-2z^{15}+3z^{10}+z^{5}\\ f_{1}+g_{1}&z^{5}f_{1}+2z^{10}+3z^{5}+2\end{pmatrix}$ $\begin{pmatrix}z^{5}&-z^{20}f_{1}-u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{25}u-z^{15}&z^{5}\end{pmatrix}$ $(q^{55},q^{55},q^{55},q^{75})$ 96 0
$\frac{50}{60}$ $z^{20}$ $\begin{pmatrix}f_{1}-u-1+2z^{10}&-f_{1}+2u+1\\ f_{1}-u+2z^{10}&-f_{1}+u-z^{20}\end{pmatrix}$ $\begin{pmatrix}z^{20}&-f_{1}+2u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ f_{1}-u+1&z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},1)$ 96 360
@@
$\frac{34}{60}$ $z^{4}$ $\begin{pmatrix}-z^{8}-z^{6}&z^{18}f_{1}+2z^{12}+4\\ z^{14}-z^{50}v+2z^{10}+u&u-z^{50}v+1\end{pmatrix}$ $\begin{pmatrix}z^{4}&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ \frac{1}{2}\left(z^{4}f_{1}+3z^{24}-z^{12}+z^{6}\right)&z^{4}\end{pmatrix}$ $(z^{28},z^{28},z^{28},z^{36})$ 60 0
$\frac{46}{60}$ $z^{16}$ $\begin{pmatrix}z^{14}-z^{4}+z^{2}&f_{2}\\ -v+z^{28}v+z^{10}-z^{8}&\left(z^{50}+z^{28}\right)v-u\end{pmatrix}$ $\begin{pmatrix}z^{16}&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ \frac{1}{2}\left(z^{22}+2z^{16}u+z^{10}-3z^{6}\right)&z^{16}\end{pmatrix}$ $(z^{22},z^{22},z^{22},z^{54})$ 60 0
@@ Appendix B Middle Convolution Tables of Nice 4-Tuples
A $\frac{1}{2}$ 1-1 $\begin{pmatrix}u&-u+1\\ u-1&-u+2\end{pmatrix}$ $\begin{pmatrix}0&1\\ -1&2\end{pmatrix}$ $\begin{pmatrix}1&u+2\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&1\end{pmatrix}$ $(1,1,1,1,1)$ - 0
Table 38: Here $q=\zeta_{5}$ and $u=q^{3}+q^{2}$
@@
Table 39: Here $z=\zeta_{3}$ and $u=\zeta_{12}$
@@
A $\frac{5}{6}$ $z^{5}$ $\begin{pmatrix}-f_{2}+1&\frac{1}{4}g_{1}\\ g_{2}&z^{5\>}+f_{2}\end{pmatrix}$ $\begin{pmatrix}f_{1}+2&\frac{1}{2}f_{2}\\ g_{3}&g_{4}\end{pmatrix}$ $\begin{pmatrix}z^{5}&\frac{1}{4}g_{5}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ g_{6}&z^{5}\end{pmatrix}$ $(z^{5},z^{5},z^{5},z^{5},z^{10})$ - 360
@@
$\frac{5}{6}$ $z^{4}+1$ $\begin{pmatrix}1&1\\ 0&z^{8}\end{pmatrix}$ $\begin{pmatrix}z^{4}+1&z^{4}\\ -2&-1\end{pmatrix}$ $\begin{pmatrix}z^{8}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 2z^{8}&z^{8}\end{pmatrix}$ $(z^{8},z^{8},z^{8},z^{8},-z^{4})$ 45 72
@@
$\frac{7}{10}$ $z^{12}$ $\begin{pmatrix}z^{12}-v&-z^{24}\\ -z^{18}+v&z^{6}\end{pmatrix}$ $\begin{pmatrix}-z^{24}+z^{12}&z^{12}\\ z^{24}-z^{6}v&z^{24}+1\end{pmatrix}$ $\begin{pmatrix}z^{12}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{24}+z^{12}-1&z^{12}\end{pmatrix}$ $(z^{24},z^{24},z^{24},z^{24},z^{24})$ 50 600
@@
$\frac{1}{6}$ $z^{10}$ $\begin{pmatrix}-z^{20}-z^{16}&-z^{16}-z^{10}\\ z^{16}-1&z^{16}\end{pmatrix}$ $\begin{pmatrix}-z^{10}&0\\ z^{10}&1\end{pmatrix}$ $\begin{pmatrix}-z^{10}&u-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{10}&-z^{10}\end{pmatrix}$ $(z^{10},z^{10},z^{10},z^{10},-z^{20})$ 90 360
$\frac{5}{6}$ $z^{10}+1$ $\begin{pmatrix}y+z^{20}&y+z^{10}-2\\ -y&-y+1\end{pmatrix}$ $\begin{pmatrix}z^{20}&0\\ -z^{20}&1\end{pmatrix}$ $\begin{pmatrix}z^{20}&u-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{20}&z^{20}\end{pmatrix}$ $(z^{20},z^{20},z^{20},z^{20},-z^{10})$ 90 360
@@
$\frac{9}{10}$ $z^{24}$ $\begin{pmatrix}1&-u-1\\ 0&z^{24}\end{pmatrix}$ $\begin{pmatrix}z^{24}+1&z^{6}v\\ u&0\end{pmatrix}$ $\begin{pmatrix}z^{24}&u+1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ u+z^{12}&z^{24}\end{pmatrix}$ $(z^{18},z^{18},z^{18},z^{18},z^{48})$ 50 600
@@
$\frac{5}{6}$ $z^{2}+1$ $\begin{pmatrix}0&-1\\ -1&0\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{2}\end{pmatrix}$ $\begin{pmatrix}z^{2}&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{2}\end{pmatrix}$ $(-z^{3},z^{5},z^{5},z^{5},1)$ 40 72
@@
$\frac{5}{6}$ $z^{2}+1$ $\begin{pmatrix}1&-w\\ 0&-1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-w^{2}\end{pmatrix}$ $\begin{pmatrix}-1&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-w\end{pmatrix}$ $(z^{3},z^{5},z^{3},z,1)$ 0
@@
$\frac{7}{12}$ z-z $\begin{pmatrix}z^{3}&0\\ -z^{11}&1\end{pmatrix}$ $\begin{pmatrix}1&z^{4}\\ 0&z^{3}\end{pmatrix}$ $\begin{pmatrix}z^{3}&-z^{4}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{11}&z^{3}\end{pmatrix}$ $(q^{9},q^{9},q^{9},q^{9},-1)$ 20 96
@@ Appendix C Middle Convolution Tables for Nice 5-Tuples
A $\frac{5}{6}$ w-w $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}0&w+1\\ -w&0\end{pmatrix}$ $\begin{pmatrix}w+1&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $(z^{11},z^{11},-z^{3},z^{11},z^{11},z^{7})$ - 72
B $\frac{5}{6}$ w-w $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $\begin{pmatrix}0&-w-1\\ w&0\end{pmatrix}$ $\begin{pmatrix}w+1&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&w+1\end{pmatrix}$ $(z^{3},z^{11},z^{3},z^{11},z^{11},-z^{3})$ - 72